# Archilles’ Heel in Semi-open LLMs: Hiding Bottom against Recovery Attacks

### v1

@@ Abstract
Closed-source large language models deliver strong performance but have limited downstream customizability. Semi-open models, combining both closed-source and public layers, were introduced to improve customizability. However, parameters in the closed-source layers are found vulnerable to recovery attacks. In this paper, we explore the design of semi-open models with fewer closed-source layers, aiming to increase customizability while ensuring resilience to recovery attacks. We analyze the contribution of closed-source layer to the overall resilience and theoretically prove that in a deep transformer-based model, there exists a transition layer such that even small recovery errors in layers before this layer can lead to recovery failure. Building on this, we propose SCARA [^1] Code is available at: https://github.com/OTTO-OTO/SCARA-Semi-Open, a novel approach that keeps only a few bottom layers as closed-source. SCARA employs a fine-tuning-free metric to estimate the maximum number of layers that can be publicly accessible for customization. We apply it to five models (1.3B to 70B parameters) to construct semi-open models, validating their customizability on six downstream tasks and assessing their resilience against various recovery attacks on sixteen benchmarks. We compare SCARA to baselines and observe that it generally improves downstream customization performance and offers similar resilience with over 10 times fewer closed-source parameters. We empirically investigate the existence of transition layers, analyze the effectiveness of our scheme and finally discuss its limitations.
@@ 1 Introduction
Figure 1: Semi-open vs. fully-closed model.
Vendors of Large Language Models (LLMs) have recently launched several closed-source models with impressive capabilities, serving diverse user needs across different scenarios (Minaee et al., 2024; Zhao et al., 2023). Access to closed-source models is typically through black-box APIs, such as GPT-4o and Claude 3.5. These APIs hide model weights and internal structures, restricting downstream customizability. To address this, semi-open LLMs have emerged, offering more flexibility while still hiding certain infrastructure details (Eiras et al., 2024; White et al., 2024). As shown in Figure 1, these semi-open models keep some modules closed-source but allow access to others. This enables users to interact with closed-source modules via embedding APIs and fine-tune the open modules for downstream tasks.
Open-sourcing more parameters and structure details apparently enhances downstream customizability. However, Zanella-Beguelin et al. (2021) showed that semi-open LLMs with only a few closed-source parameters are vulnerable to model recovery attacks. Recovery attackers query the closed-source module and then train a new module that imitates its functionality. This can lead to the full replication and theft of closed-source modules (Solaiman, 2023). Although the community has extensively studied recovery attacks against closed-source models (Chen et al., 2023; Jiang et al., 2023b), defending against recovery attacks in the semi-open setting is an uncharted area. Recovery attackers targeting fully closed-source models seek to fine-tune a new model that precisely replicates the closed-source model (Tamber et al., 2024; Dubiński et al., 2024). In contrast, attackers in semi-open settings are not required to exactly replicate the closed-source module. Instead, they can fine-tune the closed-source module alongside the public module to reconstruct the overall functionality. While open-sourcing more layers enhances downstream flexibility, it also facilitates easier replication.
Beyond the closed-source amount, the specific sections concealed are vital for defending against recovery attacks. Shen et al. (2023) suggested concealing several top layers and keeping the bottom public. However, the benefits of hiding bottom layers (near the input) versus top layers (near the output) are still unclear. Therefore, we examine the impact of each layer on resilience and theoretically identify a transition layer. Any recovery error in bottom layers before this transition layer leads to a high probability of recovery failure. In contrast, errors in later layers have limited impact. This finding suggests that keeping the bottom layers closed-sourced offers better protection against recovery attacks than the top layers, even when the same number of layers is hidden.
In this paper, we introduce SCARA, a selective closed-sourcing method for designing semi-open models that balances customizability with resilience against recovery attacks. Building on our theoretical findings, SCARA selectively hides a few bottom layers. It determines the minimal number of closed-source layers using a recovery difficulty score, a metric that estimates recovery performance without requiring fine-tuning. This score is based on the initial average recovery loss during the attack. SCARA identifies the optimal closed-source strategy by selecting the layers with metric corresponding to the worst recovery performance. Consequently, models designed by SCARA retain most layers as publicly accessible, achieving customizability comparable to fully open-source models while remaining resilient to recovery attacks. Our main contributions are as follows:
  • We theoretically demonstrate the existence of a transition layer in LLMs. We prove that small recovery errors in bottom decoder layers before this layer can lead to recovery failure with high probability, whereas errors in later layers have a limited impact. (see Section 4.1)
  • We introduce SCARA, a selective closed-source approach that conceals a few bottom decoder layers to enhance customizability while maintaining resilience. Specifically, we propose a metric that does not require fine-tuning, but correlates with the recovered performance under attacks, enabling us to approximately find the minimal number of hidden layers. (see Section 4.2)
  • We compare our approach with two baselines across five models (1.3B to 70B parameters), assessing customizability on six tasks and resilience against three recovery strategies across sixteen benchmarks in six domains. Experiments show that our method significantly improves downstream performance while maintaining comparable resilience against recovery attacks, with over 10 times fewer closed-source parameters than the baselines. For example, the semi-open Llama2-70B produced by our method hides only 2.5% of the parameters but achieves a 30% higher downstream performance score than the baselines in the Financial domain. We also observe a performance improvement of over 40% on Mistral-7B. Additionally, our method maintains similar resilience against recovery attacks compared to both baselines. (see Section 5.2)
  • We empirically investigate the presence of transition layers and the correlation between our metric and the recovered performance of each closesd-source combination. We conclude by analyzing the hyper-parameter sensitivity and discuss the limitations of our approach. (see Section 5.4)
@@ 2 Related Works
Model Customization. Vendors have introduced three main strategies for model customization, each with distinct trade-offs. First, fine-tuning APIs allow users to customize fully closed-source models (e.g., La Plateforme, Azure AI Services), while restricting access (Finlayson et al., 2024). Second, embedding models provide richer customization by allowing users to select and modify subsequent structures (Sarıtaş et al., 2024; Lee et al., 2024). However, the absence of joint pre-training can degrade performance (Nussbaum et al., 2024) and expose the closed-sourced models to recovery attacks (Caspari et al., 2024; Tamber et al., 2024). Third, open-source models enable full customization flexibility through free fine-tuning, yet pose challenges to model control and usage supervision (Bommasani et al., 2022; Roumeliotis et al., 2023).
Model Recovery Attacks. Prior attacks (Tramèr et al., 2016; Krishna et al., 2020; Dziedzic et al., 2023a) attempt to recover the functionality of fully closed-source models through API queries.  Carlini et al. (2024) advanced these by entirely extracting the embedding projection layer and hidden dimension size. Recently, various defenses against fully closed-source model recovery attacks have been proposed (Jiang et al., 2024), including malicious queries detection (Shang et al., 2024), watermarking (Zhang et al., 2021), fingerprinting (Guan et al., 2022), etc. These methods do not directly apply to the semi-open settings with only partial model information.
Semi-Open Model. Previous studies (Lin et al., 2024; Chen et al., 2024; Qiao & Zhou, 2023) explore opening bottom layers of models for user customization, while keeping later layers closed-sourced to maintain vendor control. For example, Shen et al. (2023) introduced SAP, which open-sources the first six transformer layers but limits customization options. Meanwhile, Dubiński et al. (2024); Dziedzic et al. (2023b) proposed a semi-open approach where encoder models are offered as APIs, allowing users to customize task-specific subsequent modules. However, (Liu et al., 2022) and Sha et al. (2023) showed that these encoder models are still vulnerable to recovery attacks.
@@ 3 Preliminaries
3.1 Security Threat: Semi-open Model Recovery
Semi-open LLMs. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ denote the input data matrix, where each row corresponds to a $d$-dimensional feature vector representing a single token. Let $f:\mathbb{R}^{n\times d}\rightarrow\mathcal{Y}$ denote a victim model, capable of processing the feature matrix $\mathbf{X}$ and producing an element in the set $\mathcal{Y}$ as output. Modern LLMs typically adopt a multi-layer architecture to capture complex patterns in the input data. Specifically, $f$ is a composition of multiple decoder layers, i.e., $f(\mathbf{X};\bm{\theta})=\varphi_{L}\circ...\circ\varphi_{1}(\mathbf{X})$. All decoder layers $\varphi_{1},...,\varphi_{L}$ share the same architecture but each layer is equipped with distinct parameters. The parameters of all layers are denoted by the vector $\bm{\theta}$. We consider a semi-open setting, in line with  Zanella-Beguelin et al. (2021) and Xu et al. (2021), where certain layers of the LLM are closed-sourced while others remain public. Let the closed-sourced set $I\subseteq\{1,\ldots,L\}\triangleq[L]$ denote the index set of hidden layers, while its complement $I^{c}$ contains the public layer indices.
Figure 2: Workflow of semi-open model recovery attack
Semi-open Model Recovery Attack. The semi-open model recovery attack is one of the high-fidelity attacks (Carlini et al., 2024) that aims to recover the general functionality of the victim model. As illustrated in Figure 2, the adversary is able to interact with the victim LLM and possesses access to the architecture and parameters of its public layers. The adversary aims to emulate the behavior of the victim LLM by learning the architecture and parameters associated with the closed-source layers. This task is accomplished through a structured procedure comprising three main steps: (1) Attack Dataset Construction: The adversary begins by querying the victim model, thereby gathering an attack dataset $\mathcal{D}$ containing samples representing the capabilities of the victim LLM. (2) Parameter Initialization: Next, the adversary randomly initializes the parameters associated with the closed-sourced layers, setting the stage for subsequent model fine-tuning. (3) Model Fine-Tuning: Leveraging the dataset $\mathcal{D}$ obtained in the first step, the adversary fine-tunes the parameters of the entire model, iteratively adjusting the parameters to better align with the observed behavior of the victim LLM. Let $\bm{\theta}_{\text{FT}}(I,\mathcal{D})$ denote the recovered parameters under the attack dataset $\mathcal{D}$ and closed-sourced set $I$. Prior work has shown that this approach enables the adversary to replicate the general functionality of the semi-open LLMs, despite limited access.
@@ 3.2 Problem Formulation
[…]
$R(I)=\frac{\mathbb{E}[s(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y)]}{\mathbb{E}[s(f(\mathbf{X};\bm{\theta}),Y)]}.$ (1)
@@
Here, $\mathbb{E}$ in the numerator reflects the expectation computed over random samples $(\mathbf{X},Y)$ drawn from $\mathbb{P}_{\mathbf{X}\times Y}$, the random attack set $\mathcal{D}$, and the random initialization of parameters within the closed-source layers during fine-tuning. Conversely, the term $\mathbb{E}$ in the denominator solely considers the expectation over random samples. With this definition, the term $R([L])$ denotes the recovery ratio of the recovered model under a fully-closed approach, where $[L]=\{1,...,L\}$. Hence, the following question arises.
Given $\varepsilon>0$, what is the smallest closed-source set $I$ for which $R(I)\leq(1+\varepsilon)R([L])$?
This question essentially asks whether it is feasible to identify a minimal closed-source index set $I$, such that, under this closed-source strategy, the resulting recovered model exhibits similarity to the model recovered under fully-closed approach. In other words, the recovery score does not surpass that of fully-closed approach by more than a factor of $(1+\varepsilon)$.
4 Methodology
@@
Figure 3: Customizability and resilience comparison in Llama2-70B. Higher scores indicate better customizability in Fig. (a) and weaker resilience in Fig. (b). Details can be found in Appendix C.1
In this section, we investigate how each layer affects customizability and resilience against recovery attacks. We begin with an experiment involving two semi-open Llama2-70B models, each with either the first two (Semi-Open-1) or the last two (Semi-Open-2) decoder layers closed-sourced. We compare their customization performance and recovered performance under the recovery attack. Figure 3 (a) and (b) show that although two semi-open models perform similarly on six downstream tasks, closed-sourcing the first two layers offers significantly greater resilience than the last two. Moreover, we compare the Semi-Open-1 model to the fully-closed model and observe that this semi-open model can achieve better customizability and comparable resilience at the same time. Therefore, we conjecture that, with a sufficient number of closed-sourced layers before a certain transition layer, a semi-open model can simultaneously achieve great customizability on downstream tasks and strong resilience against recovery attacks. In this section, we first present a theoretical result showing the existence of transition layers and then introduce our selective closed-sourcing approach.
4.1 Resilience Transition Layer in Infinitely Deep Transformers

### v2

# Position: On-Premises LLM Deployment Demands a Middle Path: Preserving Privacy Without Sacrificing Model Confidentiality

@@ Abstract
Current LLM customization typically relies on two deployment strategies: closed-source APIs, which require users to upload private data to external servers, and open-weight models, which allow local fine-tuning but pose misuse risks. In this position paper, we argue that (1) deploying closed-source LLMs within user-controlled infrastructure (on-premises deployment) enhances data privacy and mitigates misuse risks, and (2) a well-designed on-premises deployment must ensure model confidentiality—by preventing model theft—and offer privacy-preserving customization. Prior research on small models has explored securing only the output layer within hardware-secured devices to balance confidentiality and customization efficiency. However, we show that this approach is insufficient for defending large-scale LLMs against distillation attacks. We therefore introduce a semi-open deployment framework that secures only a few, carefully chosen layers, achieving distillation resistance comparable to fully secured models while preserving fine-tuning flexibility. Through extensive experiments, we show that securing bottom layers significantly reduces functional extraction risks. Our findings demonstrate that privacy and confidentiality can coexist, paving the way for secure on-premises AI deployment that balances usability and protection.
Machine Learning, Secure on-premises LLM Deployment
@@ 1 Introduction
Vendors of Large Language Models (LLMs) have introduced advanced models with remarkable capabilities to address diverse user needs (Minaee et al., 2024; Zhao et al., 2023). To enable customization and drive industry progress, vendors typically adopt two approaches. Closed-source vendors, like OpenAI, provide fine-tuning APIs that allow users to upload data to customize proprietary models such as GPT-4. In contrast, companies like Meta offer open-weight models, such as Llama3 (Dubey et al., 2024), which users can adapt within their own infrastructure, ensuring greater flexibility and control.
However, both approaches present significant limitations for privacy-sensitive users, such as financial institutions, healthcare organizations, and government agencies, which prioritize data security. Fine-tuning APIs from closed-source vendors offer encryption for sensitive data and comply with strict privacy regulations (Dong et al., 2023; Pang et al., 2024). However, their security fundamentally depends on vendor trust, raising ethical concerns and exposing users to potential data breaches (Sun et al., 2023).
Alternatively, fine-tuning open-weight models within private infrastructure—commonly referred to as on-premises deployment, where all data and model adaptation occur locally—ensures data protection and customization (Nevo et al., 2024). However, full disclosure of model architectures and weights heightens the risk of misuse by malicious actors, enabling misinformation generation, automated cyberattacks, and security bypasses (Hendrycks et al., 2023). These risks discourage vendors from releasing state-of-the-art (SOTA) models as open-weight, as uncontrolled distribution could lead to widespread harm. Beyond security concerns, maintaining high-quality open-weight models imposes substantial computational and financial costs (Wolfe et al., 2024), further disincentivizing full disclosure.
In this position paper, we argue that deploying closed-source SOTA models on-premises with restricted access to authorized users offers a middle path to mitigate privacy leakage and model misuse risks. However, prior research has shown that hardware-based extraction attacks can recover model parameters and architectures directly from CPUs and memory (Hu et al., 2020) within local infrastructures, enabling unauthorized users to steal proprietary model information from the deployment server. This underscores the need for stronger security measures. Therefore, we further argue that a well-designed on-premises deployment strategy must fully prevent model theft while preserving data privacy, ensuring no reliance on untrusted servers.
Our work provides an affirmative step toward addressing the fundamental challenge of balancing security and customization in on-premises LLM deployment. Existing approaches have sought to protect proprietary models using Trusted Execution Environments (TEEs) (Narra et al., 2019), but fully enclosing models within TEEs incurs prohibitive computational costs, limiting their practicality (Nayan et al., 2024). Prior research has attempted to mitigate this trade-off by securing only critical layers, such as the output layer, while leaving the remaining layers exposed for fine-tuning (Zhang et al., 2024a; Mo et al., 2020). However, our findings align with prior work showing that securing only the output layer is insufficient against distillation attacks (Zanella-Beguelin et al., 2021). Extending these attacks to Llama2-70B, we confirm that this vulnerability persists at scale, enabling near-complete functionality distillation in six domains, as shown in Figure 1. These vulnerabilities raise skepticism about whether model confidentiality and customization can truly coexist in on-premises deployment, highlighting the need for a security paradigm beyond final-layer protection. Without a viable solution, vendors risk intellectual property theft, while users must compromise data privacy when relying on external servers.
Figure 1: Semi-open Deployment.
Our work suggests that this dilemma can be resolved. We propose a theoretically inspired semi-open deployment strategy that secures only a few, carefully chosen layers, balancing security and customization while mitigating distillation risks. This enables secure on-premises AI adoption in privacy-sensitive industries such as healthcare, finance, and government. More broadly, it highlights the need for hybrid security solutions that balance usability and protection, lowering barriers to responsible AI deployment.
By demonstrating that confidentiality and customization can coexist, our work shifts the narrative from skepticism to opportunity, paving the way for AI-driven innovation in regulated sectors. Future research should explore optimal layer selection, adaptive security mechanisms, and distillation-resistant architectures to further enhance model confidentiality. We hope this work inspires the AI community to develop scalable solutions that ensure AI remains both secure and widely accessible. We summarize our contribution as follows.
  • Semi-open on-premises LLM deployment. We propose a semi-open deployment framework that secures a single, carefully chosen layer, achieving distillation resistance comparable to full encryption while preserving customization flexibility comparable to full parameter fine-tuning.
  • Layer-wise security transitions. We provide the first theoretical result identifying a transition layer that resists distillation attacks, revealing fundamental differences in security across model layers.
  • Comprehensive empirical evaluations. We have conducted extensive empirical evaluations across five models (Llama2 (70B,7B), Mistral-B, Phi-2 (2.7B), and Phi-1.5 (1.3B) ) under three distillation strategies with a 51k attack set, comparing its performance with three baselines (DarkneTZ, SAP-DP, Fully-secured).
  • Discussion of alternative positions. We objectively examine alternative positions, and highlight their benefits and limitations, reinforcing the need for a new framework that balances security, privacy, and customization. Our discussion aims to stimulate further advancements and foster new solutions in this evolving field.
@@ 2 Related Works
2.1 Data Privacy Risks
Using LLM services for customization introduces significant data privacy risks, as user data is exposed during transmission, storage, and processing (Li et al., 2024b). To address this, prior research has explored Homomorphic Encryption (Lee et al., 2022), which allows computations on encrypted data, and Differential Privacy (Wei et al., 2021), which injects noise to prevent sensitive data memorization. However, these methods suffer from high computational overhead (Zhou et al., 2024) and limited protection coverage (Wei et al., 2020). A more effective approach is on-premises deployment, which ensures data remains under user control (Nevo et al., 2024). While this mitigates privacy concerns, it shifts the risk to LLM vendors, who lose control over model usage and face heightened risks of theft and misuse (Hendrycks et al., 2023).
2.2 Model Theft Risks
Deploying LLMs on user-controlled infrastructure increases exposure to model extraction threats (Atli et al., 2020). Attackers can recover model parameters through hardware-based techniques, including side-channel attacks targeting GPUs (Nayan et al., 2024) and bus monitoring attacks (Hu et al., 2020). In addition to direct parameter extraction, distillation attacks allow adversaries to replicate model functionality using only black-box access, a process known as functional distillation (Nevo et al., 2024; Xu et al., 2024; Ezzeddine et al., 2024). While distillation attacks have been extensively studied in smaller models, such as CNNs (Orekondy et al., 2018), BERT (Sanh et al., 2020; Zanella-Beguelin et al., 2021), and ReLU-based models (Canales-Martínez et al., 2024; Jagielski et al., 2020), their effectiveness against large-scale LLMs remains an open question. Our work extends these attacks to Llama2-70B and demonstrates that securing only the output layer remains insufficient to prevent near-complete functionality replication.
2.3 Traditional Approaches
Balancing data privacy and model security in on-premises deployment requires novel solutions. One direction is privacy-preserving federated deployment, where a model is split between user-controlled infrastructure and vendor-managed servers (Shen et al., 2023; Huang et al., 2024). Users fine-tune local layers while the vendor controls deeper layers, with privacy-preserving techniques like differential privacy (Zhou et al., 2024) protecting sensitive data. However, industries with strict data regulations, such as banking, healthcare, and government, often require full on-premises deployment (Schillaci, 2024; Guerra-Manzanares et al., 2023), ensuring models remain within their infrastructure while allowing customization.
To secure models in on-premises settings, research has explored hardware-based protections such as Arm TrustZone (Pinto & Santos, 2019) and secure execution environments (Zhang et al., 2024b; Li et al., 2024a). These approaches isolate computations to prevent unauthorized extraction but impose high resource demands and restrict customization (Mo et al., 2020). A more flexible alternative is layer-wise security, where only critical layers are protected while others remain exposed (Lin et al., 2024; Chen et al., 2024; Zhang et al., 2024a). Studies suggest different strategies, including securing shallow layers (Elgamal & Nahrstedt, 2020), intermediate layers (Shen et al., 2022), or the output layer (Huang et al., 2024). However, most research has focused on smaller models, leaving the effectiveness of different security placements in LLMs unclear.
Our work investigates the layer-wise security in LLMs, showing that securing bottom layers better resists distillation. This suggests that protecting a single, well-chosen layer can defend against attacks while maximizing fine-tuning flexibility, enabling more secure on-premises deployment.
@@ 3 Preliminaries
3.1 Security Threat: Model Distillation
Semi-open LLMs. Let $f:\mathbb{R}^{n\times d}\rightarrow\mathcal{Y}$ denote a victim model that processes the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ to produce an output in set $\mathcal{Y}$. Modern LLMs typically adopt a multi-layer architecture to capture complex patterns in the input data. Specifically, $f(\mathbf{X};\bm{\theta})=\varphi_{L}\circ...\circ\varphi_{1}(\mathbf{X})$ is a composition of multiple decoder layers $\varphi_{1},...,\varphi_{L}$, each shares the same architecture but with distinct parameters. The parameters of all layers are denoted by the vector $\bm{\theta}$. We consider a semi-open setting, in line with (Zanella-Beguelin et al., 2021) and (Xu et al., 2021), where certain layers of the LLM are secured while others remain public.
Figure 2: Workflow of model distillation attack
Adversary’s Objective. The adversary aims to replicate the functionality of a semi-open victim LLM, partially secured in a protected environment, by building a replacement model. The agreement between the victim and replicated models is measured through accuracy and fidelity on a specified testing dataset.
Adversary’s Knowledge. It is assumed that the adversary knows the architectures of both secured and unsecured modules, as prior work (Gou et al., 2021; Boix-Adsera, 2024) has shown that using the same architecture as the secured module significantly improves the effectiveness of distillation attacks. However, the adversary knows only the parameters of the unsecured module, while those of the secured module remain unknown.
Adversary’s Capability. The adversary is capable of querying the semi-open model, obtaining both the semantic output produced by the complete model and the representation vector generated by the secured module. Utilizing this information, the adversary constructs a distillation attack dataset denoted as $\mathcal{D}$. Since the adversary knows the architecture of the secured module, the adversary next replaces the secured module with a randomly initialized module of the same architecture. Using the constructed set $\mathcal{D}$, the adversary employs three distinct supervised distillation strategies to replicate the functionality of the secured module: (1) FT-all: Fine-tunes both the replacement and unsecured modules using output of the entire model as training labels. (2) FT-closed: Fine-tunes only the replacement model using output of the entire model, keeping the unsecured module fixed. (3) SEM (Tamber et al., 2024): Fine-tunes the replacement model using outputs from the secured module without involving the unsecured component.
@@ 3.2 Problem Formulation
In this paper, we analyze the performance of a large language model under a defined distribution $\mathbb{P}_{\mathbf{X}\times Y}$, describing the relationship between input matrix $\mathbf{X}$ and label $Y$. We assume the victim LLM $f(\mathbf{X};\bm{\theta})$ performs well on this distribution, and the attack set $\mathcal{D}$ comprises samples drawn from $\mathbb{P}_{\mathbf{X}\times Y}$. To evaluate agreement between the distilled LLM and ground-truth labels, we use a scoring function $s:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$. Secured layers are indexed by $I\subseteq[L]=\{1,\dots,L\}$. Let $\bm{\theta}_{\text{dist}}(I,\mathcal{D})$ represent the parameter vector of the distilled replica of a victim model, where layers indexed by $I$ are secured, and adversaries utilize the attack set $\mathcal{D}$ to replicate its functionality. For each secured set $I$, we define the "Distillation Ratio" $R(I)$, which quantifies how well the distilled model $\bm{\theta}_{\text{dist}}(I,\mathcal{D})$ replicates the behavior of $f(\mathbf{X};\bm{\theta})$, expressed as
$R(I)=\frac{\mathbb{E}[s(f(\mathbf{X};\bm{\theta}_{\text{dist}}(I,\mathcal{D})),Y)]}{\mathbb{E}[s(f(\mathbf{X};\bm{\theta}),Y)]}.$ (1)
@@
Here, $\mathbb{E}$ in the numerator reflects the expectation computed over random samples $(\mathbf{X},Y)$ drawn from $\mathbb{P}_{\mathbf{X}\times Y}$, the random attack set $\mathcal{D}$, and the random initialization of parameters within the secured layers during fine-tuning. Conversely, the term $\mathbb{E}$ in the denominator solely considers the expectation over random samples. With this definition, $R([L])$ represents the distillation ratio when the entire model is secured, reflecting the highest level of security. This leads to the question:
What is the smallest secured set $\text{What is the smallest secured set }I\text{ such that }$
$\quad\quad R(I)\text{ closely approximates }R([L])?$
This question aims to identify the minimal secured set $I$ such that securing the layers indexed by $I$ achieves a level of security comparable to securing the entire model.
4 Impact of Secured Layers Placement
@@
Figure 3: Security and adaptability comparison in Llama2-70B. Lower scores indicate better security in Fig. (a) and weaker adaptability in Fig. (b). Details can be found in Appendix C.1
In this section, we investigate the impact of securing specific layers on security and customization against distillation attacks. We begin with an experiment with two semi-open deployments of Llama2-70B: one securing the bottom two decoder layers (Bottom2-Secured) and the other securing the top two decoder layers (Top2-Secured). As shown in Figure 3, both deployments achieve similar customization performance in six downstream tasks. However, securing the bottom layers provides significantly stronger security. Additionally, comparing Bottom2-Secured to fully-secured deployment reveals comparable security with improved customizability. This suggests that securing a certain number of bottom layers can effectively balance strong security against distillation attacks and high customization performance.
4.1 Security Transition in Deep Transformers